# GUARD: Guided Unlearning and Retention via Data Attribution for Large Language Models

### v1

@@ Abstract
Unlearning in large language models (LLMs) is becoming increasingly important due to regulatory compliance, copyright protection, and privacy concerns. However, a key challenge in LLM unlearning is unintended forgetting, where the removal of specific data inadvertently impairs the utility of the model and its retention of valuable, desired information. While prior work has primarily focused on architectural innovations, the influence of data-level factors on unlearning performance remains underexplored. As a result, existing methods often suffer from degraded retention when forgetting high-impact data. To address this, we propose GUARD—a novel framework for Guided Unlearning And Retention via Data attribution. At its core, GUARD introduces a lightweight proxy data attribution metric tailored for LLM unlearning, which quantifies the “alignment” between the forget and retain sets while remaining computationally efficient. Building on this, we design a novel unlearning objective that assigns adaptive, nonuniform unlearning weights to samples, inversely proportional to their proxy attribution scores. Through such a reallocation of unlearning power, GUARD mitigates unintended losses in retention. We provide rigorous theoretical guarantees that GUARD significantly enhances retention while maintaining forgetting metrics comparable to prior methods. Extensive experiments on the TOFU benchmark across multiple LLM architectures demonstrate that GUARD substantially improves utility preservation while ensuring effective unlearning. Notably, GUARD reduces utility sacrifice on the Retain Set by up to 194.92% in terms of Truth Ratio when forgetting 10% of the training data.
@@ 1 Introduction
Although LLMs have demonstrated remarkable domain transferability and achieved state-of-the-art performance across a wide range of natural language processing tasks [33, 47, 43, 48, 24], their deployment has raised serious privacy and security concerns. LLMs are typically trained on massive corpora of data that may contain sensitive, private, or proprietary information, which increases the risk of unintended data leakage and misuse [21, 39, 22, 49] (e.g., some recent studies have shown that LLMs can memorize and regenerate specific pieces of sensitive content, such as Social Security Numbers (SSNs) [3, 11]). These challenges have prompted a growing interest in developing mechanisms for the selective removal — or unlearning — of specific information from LLMs.
@@
Machine Unlearning in LLMs provides a practical alternative to avoid model retraining when removing specific information in compliance with privacy and regulatory requirements [52, 18, 4, 30, 26]. Despite many advances in the domains of computational efficiency and key forget parameter identification [50, 29, 53, 20, 40, 6, 46, 15], fine-tuning-based LLM unlearning methods frequently suffer from excessive forgetting and retention degradation [54], compromising the utility of the retained model knowledge. Similarly, inference-time LLM unlearning techniques [35, 12, 41, 27, 14], which avoid model updates and are hence computationally attractive, do not guarantee actual forgetting, as sensitive information remains embedded within the unchanged model parameters. Related work for both lines is in Appendix A.1. More broadly, the current limitations of LLM unlearning methods include one or more of the following: a) degraded retention, which refers to the fact that unlearning can severely degrade performance on the retained data [54, 53, 20]; b) computational inefficiency, which arises when requiring auxiliary LLMs for guidance or evaluation [14]; c) privacy leakage, which results in failure to guarantee acceptable removal of sensitive information [27, 35, 12]; and, d) lack of theoretical performance guarantees. To address the above problems, we propose a novel unlearning method that ensures increased utility, effective forgetting, improved computational efficiency, and formal theoretical guarantees for both retention and forgetting.
@@
Increasing utility and effective forgetting. Existing unlearning methods use unprincipled allocation of model unlearning weights across forget samples. For example, prior fine-tuning-based approaches typically construct the unlearning objective by averaging per-sample losses with retention-unrelated weights that are often uniform or determined solely by the forget sample itself, without considering the influence of the sample on model performance over the retained set. Neglecting the attributional impact of a sample leads to unintentional forgetting of useful knowledge and consequent utility loss: for example, unlearning a training instance from a particular class can disproportionately degrade performance on that same class, compared to unlearning unrelated instances [17]. Our solution involves a novel data-attribution-driven unlearning framework. To address the limitations of retention-unaware unlearning, we propose GUARD (Guided Unlearning And Retention via Data attribution), a novel unlearning framework that explicitly incorporates data attribution metrics into the unlearning objective; GUARD dynamically adjusts unlearning weights based on the influence of each forget sample on the performance of the model over the retained data. More concretely, GUARD allocates unlearning weights in such a way that the data-wise loss weight is inversely proportional to the sample-wise contribution to retention, reducing the likelihood of unintentionally degrading retained knowledge. To ensure effective forgetting, we also introduce a temperature-controlled reverse unification mechanism that regulates the variance of unlearning weights across samples. This principled reallocation strategy enables GUARD to both mitigate retention degradation by down-weighting high-retention-impact samples, and to maintain effective forgetting by carefully calibrating the degree of adjustment.
@@
Improved computational efficiency. Since GUARD has to perform efficient data attribution, one has to devise appropriate methods for LLMs. Conventional data attribution methods estimate the influence of individual training samples on overall model performance by comparing model outputs with and without the inclusion of each sample [17, 9, 13]. However, these approaches typically require model retraining or Hessian-based approximations, making them computationally prohibitive for LLMs. Moreover, evaluating the influence of samples based on global performance contradicts the unlearning objective which aims to decrease the performance on the forget set and simultaneously preserve or improve it on the retain set, rather than optimize performance across the whole dataset (see Appendix A.2). Our solution involves novel attribution scores tailored to LLM unlearning. Inspired by gradient-based sample-wise representations [32, 45], we define the retention attribution score of a forget sample as the inner product between its gradient and the gradient averaged over all retain samples. This measure captures the alignment between the knowledge encoded by a forget sample and that of the retain distribution. It also avoids the computational burden of retraining or Hessian estimation inherent to prior attribution techniques. Furthermore, by explicitly separating forget and retain knowledge before computing inter-sample influence, the proposed score effectively quantifies the contribution of a sample to retention utility, and facilitates retention-aware unlearning.
@@
Theoretical Guarantees. We provide a rigorous theoretical analysis demonstrating that, compared to conventional unlearning baselines such as Gradient Ascent [50], GUARD offers the following key benefits: (1) it effectively reduces the loss on the retain set, leading to enhanced absolute retention; (2) it maintains comparable degradation on the forget set, ensuring strong absolute forgetting; and (3) it significantly reduces the retention-performance tradeoff, improving relative retention efficiency, i.e., retention preserved per unit of forgetting incurred.
@@
Empirical Validation. To empirically validate our claims, we conduct extensive experiments on the TOFU benchmark (Task of Fictitious Unlearning for LLMs) [31], a standardized evaluation suite designed specifically for LLM unlearning. Across a range of LLM architectures and evaluation protocols, GUARD consistently and substantially outperforms existing methods both with respect to absolute and relative retention, while it maintains comparable levels of forgetting. Notably, GUARD reduces utility degradation on the retain set by up to 194.92% in terms of the Truth Ratio when forgetting 10% of the training data.
@@
  • Extensive Empirical Validation. We evaluate GUARD on the TOFU benchmark and show that it consistently outperforms baselines in both absolute and relative retention, while maintaining comparable forgetting, e.g., it reduces retain-set utility loss by up to 194.92% in Truth Ratio over GA baseline when forgetting 10% of the training data.
@@ 2 LLM Unlearning Preliminaries
Data: The forget and retain sets. We define a datapoint as $(x,y)$, where $x\in\mathbb{R}^{p}$ represents the input of the LLM, while $y\in\mathbb{R}^{q}$ denotes the corresponding ground-truth label or response. The whole dataset used for fine-tuning an LLM is denoted as $D_{0}=\{(x_{i},y_{i})\}_{i\in[n_{0}]}$, where $n_{0}$ is the total number of datapoints. The forget set, denoted by $D_{f}$, consists of datapoints requested for removal, while the retain set, denoted by $D_{r}$, contains datapoints that should be kept. The index sets of these datapoints are defined as $I_{f}:=\{i\mid(x_{i},y_{i})\in D_{f}\}$ and $I_{r}:=\{j\mid(x_{j},y_{j})\in D_{r}\}$. We also let $n_{f}:=|I_{f}|$ and $n_{r}:=|I_{r}|$ represent the respective cardinalities of the sets. Clearly, $n_{f}+n_{r}=n$ and $D=D_{f}\cup D_{r}$.
@@
Standard LLM unlearning. The standard process typically involves two steps (see Fig. 2):
@@
    Finetune a pretrained LLM on the whole dataset $D_{0}$ to obtain
    $\theta_{0}=\arg\min_{\theta}{L}_{D_{0}}(\theta).$ (1)
@@
    Unlearn the finetuned LLM to remove the forget set $D_{f}$. One typical approach is Gradient Ascent (GA), which updates the model parameters as follows:
    $\theta_{GA}$ $=\theta_{0}+\eta\cdot\nabla_{\theta}{L}_{D_{f}}(\theta_{0})$
    $=\theta_{0}+\eta\cdot\sum_{i\in I_{f}}\frac{1}{n_{f}}\cdot\nabla_{\theta}l(x_{i},y_{i};\theta_{0}),\,\text{ where $\eta$ denotes the unlearningrate.}$. (2)
    Unlearning mechanisms other than Gradient Ascent are discussed in Appendix D.1.
3 Method: Guided Unlearning and Retention via Data Attribution (GUARD)
3.1 GUARD Pipeline
Figure 1: The GUARD pipeline: Step 1) Standard fine-tuning of the LLM on the target task. Step 2) Calculation of unlearning weights based on data attribution, followed by reverse unification of the scores. Step 3) GUARD unlearning removes the influence of the forget set while preserving desired knowledge.
Unlike traditional methods that assign retention-unaware unlearning weights across all data points (e.g., uniform weights as in GA, $\frac{1}{n_{f}}$; see Eq. 2), GUARD introduces a novel objective function that controllably leverages sample loss weights that are inversely correlated to data attribution, so as to model retention (Section 3.4). Additionally, GUARD uses a new data attribution scheme specifically tailored to the LLM unlearning setting (Section 3.2). The GUARD pipeline is outlined in Fig. 1, and it comprises the following steps:
  • Step 1
    Finetune the pretrained LLM: Finetuning is performed on a pretrained LLM using the full dataset, following Eq. 1.
  • Step 2
    Compute data attribution scores for retention: Estimate proxy data attribution scores for all datapoints in the forget set, as described in Section 3.2.
  • Derive retention-aware unlearning weights: Calculate unlearning weights that are inversely related to the computed attribution scores, with the degree of adjustment modulated by a temperature parameter $\tau$ (Section 3.3).
@@ 3.1 GUARD Pipeline
    Perform GUARD unlearning: Optimize the GUARD objective using the retention-aware unlearning weights, as detailed in Section 3.4.
@@ 3.2 Data Attribution for LLM Unlearning
Conventional data attribution methods estimate the influence of individual training samples by measuring the change in the model outputs when a sample is removed from the training set. These approaches typically rely on retraining or Hessian-based computations (see Appendix A.2). However, such methods are impractical in the context of LLM unlearning for two key reasons: (1) their computational cost is prohibitive at the operational scale of LLMs; and (2) the objective of unlearning differs fundamentally from standard attribution, i.e., effective unlearning necessitates different performance shifts, namely, degraded performance on the forget set ${\mathcal{D}}_{f}$ and improved performance on the retain set ${\mathcal{D}}_{r}$ (rather than a uniform unidirectional performance change over the entire dataset).
To address these challenges, we propose a proxy attribution score $a_{i}^{\text{GUARD}}$ for a forget sample $i$, defined as an inner product of gradients:
@@
where $J^{{\mathcal{D}}_{r}}_{avg}(\theta_{0})$ denotes the mean gradient over ${\mathcal{D}}_{r}$ and $J_{i}(\theta_{0})$ is the gradient of sample $i$, both evaluated for the initial model parameters $\theta_{0}$. This formulation captures the alignment between the information content of a forget and that of the retain set. Representing sample-level influence via gradients is a well-established practice, particularly in LLMs [32, 45]. In addition to being intuitive and computationally efficient, our proxy attribution score lends itself to theoretical analysis, as described in Appendix B.

### v2

@@ Abstract
Unlearning in large language models (LLMs) is becoming increasingly important due to regulatory compliance, copyright protection, and privacy concerns. However, a key challenge in LLM unlearning is unintended forgetting, where the removal of specific data inadvertently impairs the utility of the model and its retention of valuable, desired information. While prior work has primarily focused on architectural innovations, the influence of data-level factors on unlearning performance remains underexplored. As a result, existing methods often suffer from degraded retention when forgetting high-impact data. To address this problem, we propose GUARD — a novel framework for Guided Unlearning And Retention via Data attribution. At its core, GUARD introduces a lightweight proxy data attribution metric tailored for LLM unlearning, which quantifies the “alignment” between the Forget and Retain sets while remaining computationally efficient. Building on this, we design a novel unlearning objective that assigns adaptive, nonuniform unlearning weights to samples, inversely proportional to their proxy attribution scores. Through such a reallocation of unlearning power, GUARD mitigates unintended retention loss. We also provide rigorous theoretical guarantees that GUARD significantly improves retention while maintaining forgetting metrics comparable to prior methods. Extensive experiments on the TOFU and MUSE benchmarks across multiple LLM architectures demonstrate that GUARD reduces utility sacrifice on the TOFU Retain Set by up to 194.92% in terms of Truth Ratio when forgetting 10% of the training data, and improves knowledge retention on the MUSE NEWS Retain Set by 16.20%, with comparable or very moderate increases in privacy loss compared to state-of-the-art methods.
@@ 1 Introduction
Although LLMs have demonstrated remarkable domain transferability and achieved state-of-the-art performance across a wide range of natural language processing tasks Ouyang et al. (2022); Wei et al. (2022); Touvron et al. (2023); Wu et al. (2023); Liang et al. (2022), their deployment has raised serious privacy and security concerns. LLMs are typically trained on massive corpora of data that may contain sensitive, private, or proprietary information, which increases the risk of unintended data leakage and misuse Li et al. (2021); Shi et al. (2023); Li et al. (2024b); Yang et al. (2023) (for example, LLMs were shown to memorize and regenerate specific pieces of sensitive content, such as Social Security Numbers (SSNs) Carlini et al. (2021); Huang et al. (2022)). These challenges have prompted a growing interest in developing mechanisms for the selective removal — or unlearning — of specific information from LLMs.
@@
Machine Unlearning in LLMs provides a practical alternative to avoid model retraining when removing specific information, in compliance with privacy and regulatory requirements Yao et al. (2024b); Kumar et al. (2022); Chen & Yang (2023); Maini et al. (2024a); Liu & et al. (2024). Despite many advances in the domains of computational efficiency and key forget parameter identification Yao & et al. (2023); Maini & et al. (2024); Zhang & et al. (2024a); Li & et al. (2024); Shi & et al. (2024); Choi & et al. (2024); Wang et al. (2024a); Jia et al. (2024); Niu et al. (2025); Wei et al. (2025), fine-tuning-based LLM unlearning methods frequently suffer from excessive forgetting and retention degradation Zhang & et al. (2024b), compromising the utility of the retained model knowledge. Similarly, inference-time LLM unlearning techniques Pawelczyk & et al. (2023); Huang & et al. (2024); Thaker & et al. (2024); Liu et al. (2024); Ji et al. (2024), which avoid model updates and are hence computationally attractive, do not guarantee actual forgetting, as sensitive information remains embedded within the unchanged model parameters (for related works, see Appendix B.1).
More broadly, the current limitations of LLM unlearning methods include one or more of the following: a) degraded retention, which refers to the fact that unlearning can severely degrade performance on the retained data Zhang & et al. (2024b; a); Li & et al. (2024); b) computational inefficiency, which arises when requiring auxiliary LLMs for guidance or evaluation Ji et al. (2024); c) privacy leakage, which results in failure to guarantee acceptable removal of sensitive information Liu et al. (2024); Pawelczyk & et al. (2023); Huang & et al. (2024); and, d) lack of theoretical performance guarantees. To address the above problems, we propose a novel unlearning method that ensures increased utility, effective forgetting, improved computational efficiency, and formal theoretical guarantees for both retention and forgetting.
@@
Increased utility and effective forgetting. Existing unlearning methods use unprincipled allocation of model unlearning weights across forget samples. For example, prior fine-tuning-based approaches typically construct the unlearning objective by averaging per-sample losses with retention-unrelated weights that are often uniform or determined solely by the forget sample itself, without considering the influence of the sample on model performance over the retained set. Neglecting the attributional impact of a sample leads to unintentional forgetting of useful knowledge and consequent utility loss: For example, unlearning a training instance from a particular class can disproportionately degrade performance on that same class when compared to unlearning unrelated instances Koh & Liang (2017). Our solution involves a novel data-attribution-driven unlearning framework termed GUARD (Guided Unlearning And Retention via Data attribution), which explicitly incorporates data attribution metrics into the unlearning objective.
GUARD dynamically adjusts unlearning weights based on the influence of each forget sample on the performance of the model over the retained data. More precisely, GUARD allocates unlearning weights in such a way that the data-wise loss weight is inversely proportional to the sample-wise contribution to retention, reducing the likelihood of unintentionally degrading of retained knowledge. To ensure effective forgetting, we also introduce a temperature-controlled reverse unification mechanism that regulates the variance of unlearning weights across samples. This principled reallocation strategy enables GUARD to both mitigate retention degradation by down-weighting high-retention-impact samples, and to maintain effective forgetting by carefully calibrating the degree of adjustment.
@@
Improved computational efficiency. Since GUARD has to perform efficient data attribution on large LLM models, one has to redesign existing methods. Conventional data attribution methods estimate the influence of individual training samples on overall model performance by comparing model outputs with and without the inclusion of each sample Koh & Liang (2017); Ellis et al. (2022); Ingram et al. (2023). However, these approaches typically require model retraining or Hessian-based approximations, making them computationally prohibitive for LLMs. Moreover, evaluating the influence of samples based on global performance contradicts the unlearning objective which aims to decrease the performance on the Forget Set and simultaneously preserve or improve it on the Retain Set, rather than optimize performance across the whole dataset (see Appendix B.2). Our solution to this problem is a novel attribution scores tailored to LLM unlearning. Inspired by gradient-based sample-wise representations Nguyen et al. (2025); Wang et al. (2025b), we define the retention attribution score of a forget sample as the inner product between its gradient and the gradient averaged over all retain samples. This measure captures the alignment between the knowledge encoded by a forget sample and that of the retain distribution. It also avoids the computational burden of retraining or Hessian estimation inherent to prior attribution techniques. Furthermore, by explicitly separating forget and retain knowledge before computing inter-sample influence, the proposed score effectively quantifies the contribution of a sample to retention utility and facilitates retention-aware unlearning.
@@
Theoretical Guarantees. We provide a rigorous theoretical analysis demonstrating that, compared to conventional unlearning baselines such as Gradient Ascent Yao & et al. (2023), GUARD offers the following key benefits: (1) it effectively reduces the loss on the Retain Set, leading to enhanced absolute retention; (2) it maintains comparable degradation on the Forget Set, ensuring strong absolute forgetting; and (3) it significantly reduces the retention-performance tradeoff, improving relative retention efficiency, i.e., retention preserved per unit of forgetting incurred.
@@
Empirical Validation. To empirically validate our claims, we conduct extensive experiments on the TOFU Maini et al. (2024b) and MUSE Shi et al. (2024a) benchmarks, two standardized evaluation suites designed specifically for LLM unlearning. Across a range of LLM architectures and evaluation protocols, GUARD consistently and substantially outperforms existing methods both with respect to absolute and relative retention while it maintains comparable levels of forgetting. Notably, GUARD reduces utility degradation on TOFU’s Retain Set by up to 194.92% in terms of the Truth Ratio when forgetting 10% of the training data. Also, on the MUSE NEWS Retain Set, GUARD enhances knowledge retention by 16.20%.
@@
  • Extensive Empirical Validation. We conduct comprehensive experiments on the TOFU and MUSE benchmarks, covering four non-loss reweighting baselines (GA, GD, KM, PO), their GUARD-augmented counterparts (GA+GUARD, GD+GUARD, KM+GUARD, PO+GUARD), four loss reweighting baselines (NPO, SimNPO, FLAT, SimImP), as well as NPO+GUARD (we selected NPO for integration with GUARD as it demonstrated high stability and strong performance). For all these settings, GUARD consistently enhanced retention while maintaining forgetting performance comparable to baselines, demonstrating its effectiveness as a general framework applicable to diverse unlearning systems.
@@ 2 LLM Unlearning Preliminaries
Data – The Forget and Retain Sets. We define a datapoint as $(x,y)$, where $x\in\mathbb{R}^{p}$ represents the input of the LLM, while $y\in\mathbb{R}^{q}$ denotes the corresponding ground-truth label or response. The whole dataset used for fine-tuning an LLM is denoted as $D_{0}=\{(x_{i},y_{i})\}_{i\in[n_{0}]}$, where $n_{0}$ is the total number of datapoints. The Forget Set, denoted by $D_{f}$, consists of datapoints requested for removal, while the Retain Set, denoted by $D_{r}$, contains datapoints that should be kept. The index sets of these datapoints are defined as $I_{f}:=\{i\mid(x_{i},y_{i})\in D_{f}\}$ and $I_{r}:=\{j\mid(x_{j},y_{j})\in D_{r}\}$. We also let $n_{f}:=|I_{f}|$ and $n_{r}:=|I_{r}|$ represent the respective cardinalities of the sets. Clearly, $n_{f}+n_{r}=n_{0}$ and $D_{0}=D_{f}\cup D_{r}$.
@@
Standard LLM unlearning. Standard LLM unlearning refers to the process of adapting an LLM initially trained on $D_{0}$ in order to remove the influence of a designated Forget Set $D_{f}$ (see Fig. 2). One typical approach is Gradient Ascent (GA), which updates the model parameters $\theta_{0}$ as follows:
$\theta_{GA}$ $=\theta_{0}+\eta\cdot\nabla_{\theta}{L}_{D_{f}}(\theta_{0})$
$=\theta_{0}+\eta\cdot\sum_{i\in I_{f}}\frac{1}{n_{f}}\cdot\nabla_{\theta}l(x_{i},y_{i};\theta_{0}),\,\text{ where $\eta$ denotes the unlearning rate.}$ (1)
Unlearning mechanisms other than Gradient Ascent are discussed in Appendix E.1.
3 Method: GUARD
Figure 1: The GUARD pipeline: Step 1) Calculation of proxy data attribution. Step 2) Calculation of unlearning weights based on data attribution, followed by reverse unification of the scores. Step 3) Incorporation of the computed retention-aware weights for reduction of the influence of the Forget Set and preservation of desired knowledge.
3.1 GUARD Pipeline
Unlike traditional methods that assign retention-unaware unlearning weights across all data points (e.g., uniform weights as in GA, $\frac{1}{n_{f}}$; see Eq. 1), GUARD introduces a novel objective function that controllably leverages sample loss weights that are inversely correlated to data attribution scores, so as to model retention (Section 3.4). Additionally, GUARD uses a new data attribution scheme specifically tailored to the LLM unlearning setting (Section 3.2). The GUARD pipeline is outlined in Fig. 1, and it comprises the following steps:
@@ 3.1 GUARD Pipeline
    Computing the data attribution scores for retention: Estimating proxy data attribution scores for all datapoints in the Forget Set, as described in Section 3.2.
@@
    Deriving retention-aware unlearning weights: Calculating unlearning weights that are inversely related to the computed attribution scores, with the degree of adjustment modulated by a temperature parameter $\tau$ (Section 3.3).
@@
    Performing GUARD unlearning: Optimizing the GUARD objective using the retention-aware unlearning weights, as detailed in Section 3.4.
@@ 3.2 Data Attribution for LLM Unlearning
Conventional data attribution methods estimate the influence of individual training samples by measuring the change in the model outputs when a sample is removed from the training set. These approaches typically rely on retraining or Hessian-based computations (see Appendix B.2). However, such methods are impractical in the context of LLM unlearning for two key reasons: (1) their computational cost is prohibitive at the operational scale of LLMs; and (2) the objective of unlearning differs fundamentally from standard attribution, i.e., effective unlearning necessitates different performance “shifts,” namely, degraded performance on the Forget Set ${\mathcal{D}}_{f}$ and improved performance on the Retain Set ${\mathcal{D}}_{r}$ (rather than a uniform unidirectional performance change over the entire dataset). To address these challenges, we propose a proxy attribution score $a_{i}^{\text{GUARD}}$ for a Forget Sample $i$, defined as an inner product of gradients:
@@
where $J^{{\mathcal{D}}_{r}}_{avg}(\theta_{0})$ denotes the mean gradient over ${\mathcal{D}}_{r}$ and $J_{i}(\theta_{0})$ is the gradient of sample $i$, both evaluated for the initial model parameters $\theta_{0}$. This formulation captures the alignment between the information content of the Forget and the Retain Set. Representing sample-level influence via gradients is a well-established practice, particularly in LLMs Nguyen et al. (2025); Wang et al. (2025b). In addition to being intuitive and computationally efficient, our proxy attribution score lends itself to theoretical analysis, as described in Appendix C.